OpenStudy (anonymous):

Hi Can someone help me show why the following sequence isn't converging when defined as followed: x_2n-1=1/5^n , x_2n = (n)root (1/x_2n-1) .. I'll make the problem clearer in the comments.. Thanks!

OpenStudy (anonymous):

\[x _{2n-1}=1/5^n , x _{2n} = \sqrt[n]{1/2_{n-1}}\]

OpenStudy (amistre64):

what is the limit of (n)/(n-1)

OpenStudy (anonymous):

1 right?

OpenStudy (amistre64):

dunno, havent worked it out yet

OpenStudy (anonymous):

It is 1 ..

OpenStudy (anonymous):

When n approaches 0 its 1, it's 0 as n approaches infinity

OpenStudy (amistre64):

\[\sqrt[n]{\frac{1}{1/5^n}}\] \[\sqrt[n]{5^n}=5\]

OpenStudy (anonymous):

But why is it not convergent ?

OpenStudy (amistre64):

from the looks of it, it jumps between 0 and 5

OpenStudy (anonymous):

so it oscillates to infinity between 0 and 5?

OpenStudy (amistre64):

1 0.2 5 2 0.04 5 3 0.008 5 4 0.0016 5 5 0.00032 5 6 0.000064 5 7 0.0000128 5 8 0.00000256 5 9 0.000000512 5

OpenStudy (anonymous):

So in conclusion .. That can be used to prove that this sequence does not converge ?

OpenStudy (amistre64):

im not sure what 'proof' can be used x_k, for odd k goes to 0 x_k for even k maintains at 5 so there is no limiting value that it settles down at

OpenStudy (amistre64):

think of a super bouncy ball that is forever bouncing between the ceiling and the floor, where does it come to rest at?

OpenStudy (anonymous):

Eventually 0

OpenStudy (amistre64):

eventually it is either at the ceiling or on the floor, it never comes to rest

OpenStudy (anonymous):

Wait sorry no it oscillates between 0 and 5 sorry lol

OpenStudy (amistre64):

:)

OpenStudy (anonymous):

Alright I think Iunderstand

OpenStudy (anonymous):

Thanks!!!

OpenStudy (amistre64):

yep ;)